# UniGS: Unified Geometry-Aware Gaussian Splatting for Multimodal Rendering

### v1

@@ 1 Introduction
High-quality novel view synthesis (NVS) and fast multimodal 3D reconstruction remains a challenging problem in computer graphics and attracts widespread attention in fields such as autonomous driving [66, 19, 18, 78] and robotics [82, 39]. In recent years, neural radiance fields (NeRF) [55, 2, 1, 42] and 3D Gaussian Splatting (3DGS) [26, 25] have made big breakthroughs in this field. NeRF uses coordinate-based networks to store 3D scene in multi-layer perceptrons (MLPs) via ray sampling, bringing in photo-realistic reconstruction. 3DGS, another approach, uses explicit 3D Gaussian primitives and modern GPU-accelerated splatting, and this design has been shown to enable high-quality real-time NVS by many 3DGS-based frameworks [40, 29, 12, 45, 17, 65, 49, 66, 37, 38]. However, the two types of existing methods still have many issues, which mainly focus on the following aspects. NeRF requires long training times, lacks easy scene editability, and its implicit 3D information storage method renders geometric reconstruction less intuitive [79, 42]. Additionally, it performs poorly when applied to large-scale scenes [50, 53]. 3DGS and its related frameworks have more challenges: 1) Poor geometric structure and consistency: conventional image-centric 3DGS often ignores objects’ geometric structure and semantic info. It only uses Gaussian centers to represent depth, leaving out Gaussian rotation and scale attributes. This stops gradient propagation for geometric constraints (e.g. depth and normal), leading to inconsistent surface normals [40, 25, 33, 61] and misalignment between Gaussians primitives and reconstructed surfaces. 2) Inefficient densification and pruning: some 3DGS methods create too many redundant Gaussians primitives to fit RGB images well [26, 33, 25, 66]. Though this improves metrics, it slows down rendering and uses more GPU memory, losing 3DGS’s real-time edge. 3) Poor unified multimodal framework: some 3DGS-related methods use separate frameworks for multimodal rendering [32, 15], which causes repeated calculations and inconsistent 3D scene representations, failing to unify multimodal modeling. Existing studies have tried single or multiple scene modalities [40, 61, 33, 10], but there is still no unified framework that combines all these modalities while keeping storage efficient and rendering fast. This gap serves as the key motivation for our approach.
@@ 2 Related Works
Photo-Realistic Rendering and Densification. In the domain of photo-realistic reconstruction and rendering, NeRF and 3DGS represent implicit and explicit techniques, respectively. NeRF [42, 1] and its variants pioneer the use of implicit neural representations for NVS, achieving high-fidelity rendering by optimizing a continuous volumetric scene network using multi-view images. However, this approach typically requires extremely long training times. Subsequent works incorporate geometric prior structures [73, 2, 54] or efficient encoding [44] to accelerate the process. Despite these improvements, NeRF still faces inherent limitations, such as the difficulty in editing implicitly represented scenes. In contrast, 3DGS [25, 67] employs explicit Gaussian primitives as scene representations. Through compression [40, 29, 12, 45, 17] and efficient geometric structures [26, 4], several studies extend 3DGS to large-scale outdoor [65, 49, 66, 34] and urban [37, 38] environments while maintaining high-quality reconstruction. But the aforementioned research primarily focuses on RGB-based rendering and seldom incorporates explicit geometric information of the environment. Moreover, impressive efforts have been made to enhance optimization and storage efficiency. SteepestGS [58] uncovers fundamental principles of density control. DashGaussian [5] significantly accelerates 3DGS optimization. GeoTexDensifier [20] utilizes geometric information for Gaussian densification, and Pixel-GS [77] introduces pixel-level regularization for density control. These methods are regarded as post-processing techniques and lack integration within a unified framework.
@@
Geometry-Aware Rendering. Geometry-aware approaches enhance reconstruction accuracy by incorporating surfaces [61, 22, 14, 9], normals [61, 50], and relighting [13, 64, 71] capabilities, enabling more physically realistic renderings and interactions. Some studies [57, 46, 68, 50] attempt to achieve accurate geometry within the NeRF framework, and while significant progress has been made, these methods still suffer from inherent limitations that are difficult to overcome with implicit representations. 3DGS [25, 40] does not inherently excel in geometric reconstruction either, as its Gaussian primitives are independent of each other, making precise geometry recovery an open challenge. Several works unitize 3DGS to recover depth [25, 40, 56, 61], surfaces [61, 22] or meshes [14, 9], including normal estimation [61] and relighting [13, 64, 71] techniques, which can effectively reconstruct smooth object surfaces. Nevertheless, these approaches remain largely limited to small-scale objects. When extended to large environments such as outdoor or urban scenes [43, 63], the required computational time remains substantial. Some methods attempt to incorporate SDF-based surface constraints [74, 68, 59, 35, 69], inverse rendering [51, 70], or restrictions [33, 24] on Gaussian properties. However, these approaches lack rigorous analytical derivation for geometric modeling, employ relatively limited quantitative metrics, and demonstrate constrained generalizability across diverse scenarios. Consequently, the improvements achieve thus far remain marginal in terms of geometric fidelity.
@@
Semantic Rendering Frameworks. Several works attempt to integrate semantic information into both NeRF [79, 6] and 3DGS [15, 47], including techniques like Open-Vocabulary Segmentation [75, 28, 10] and Language-Driven Semantic Segmentation [48, 36, 41, 31], with promising results. Embedding semantic information facilitates downstream tasks such as scene understanding [21, 80], navigation [30, 23], and scene editing [72, 11]. For instance, SGS-SLAM [32] directly uses semantic ground truth as supervision signal and rendering output in 3DGS, though it fails to formulate a unified 3DGS map representation. Feature-GS [81] extracts semantic features from the environment using 3DGS, and SEGS-SLAM [62] enhances structural awareness with appearance embeddings. Nevertheless, such approaches exhibit limitations in reconstructing accurate environmental geometry and fail to achieve consistent reconstruction of scene geometry and semantics.
@@ 3.1 Unified Map Representation and Objectives
Objective. The objective of this framework is to render RGB color image, geometric depth, surface normals, semantic logits, and a gradient factor map (used in differentiable pruning) simultaneously within a unified rasterization pipeline. These rendered outputs are then compared against their corresponding 2D ground truth to compute respective loss, which guide the optimization of the 3D Gaussian scene through a differentiable backward process. To accelerate convergence and enhance computational efficiency, we derive the relevant analytical solutions and implement them using CUDA-accelerated gradient propagation. Finally, we can obtain a geometry consistent photo-realistic 3D Gaussian Splatting map representation.
@@ 3.3 Differentiable Pruning
Pruning. After a fixed interval of iterations $K_{p}$ ($K_{p}=3000$ in our experiments), our framework directly eliminates Gaussians with abnormal gradient factor $\|k-\mathbf{1}\|_{1}<\mathcal{T}_{k}$ , effectively pruning the set and reducing the number of Gaussians to enhance overall efficiency. After pruning operation, the $k$ value of all Gaussians will be reset to 0.9 to maintain the sustainability of updates. Please refer to our code for details.
@@ 3.4 Multimodal Rendering Framework
Architecture Overview. Our system is built upon a modified Scaffold-GS [40] architecture. For each anchor point, we maintain a foundational 3D Gaussian representation (1). During every training iteration, multiple MLPs are used to predict attribute offsets for each anchor.
@@ 24   Save $T_{\text{final}}=T$;
Method $|$ Metrics Abs.Rel↓ RMSE↓ CosSimi ↑ Abs.Rel↓ RMSE↓ CosSimi ↑ Abs.Rel↓ RMSE↓ CosSimi ↑ Abs.Rel↓ RMSE↓ CosSimi ↑
GeoGaussian [33] 0.5828 0.8006 0.3046 2.8959 1.1092 0.1861 0.1336 0.3236 0.5700 0.1082 0.2592 0.5998
3DGS [25] 1.2231 0.9183 0.2304 0.8711 0.6238 0.3201 0.2134 0.4092 0.4963 1.4422 1.0972 0.2905
Scaffold-GS [40] 2.2857 1.2102 0.2195 1.4609 0.9413 0.2854 4.0575 1.4786 0.1555 - - 0.1286
Normal-GS [61] 0.0074 0.0657 0.8755 0.0154 0.0344 0.9234 0.0243 0.0766 0.8815 0.0192 0.0589 0.8987
Ours 0.0062 0.0310 0.8880 0.0015 0.0073 0.9686 0.0136 0.0467 0.9545 0.0144 0.0465 0.9076
@@
Tile-based Forward Rasterization. We extend the original 3DGS’s tile-based rendering pipeline [25]. During rasterization, for each pixel $p=(u,v)$, we calculate the direction $\mathbf{d}$ of the camera ray $r(t)$ and then compute the intersection $t_{1}/t_{2}$ with the Gaussian projected onto this pixel. Finally, we obtain the corresponding attributes using alpha blending rendering. The multimodal forward rasterization pipeline is outlined in Alg. 2. The corresponding backpropagation pipeline is described in the supplementary materials.
Loss Function. For the RGB image loss, we employ a combination of L1 loss $\mathcal{L}_{l1}$ and SSIM [60] loss $\mathcal{L}_{\text{ssim}}$. For the normal loss $\mathcal{L}_{\text{normal}}$, we use cosine similarity, which is calculated as
@@
Implementation Details. Our framework is implemented by CUDA using the LibTorch framework [7], incorporating CUDA code for Gaussian Splatting and trained on a server with a 2.60 GHz Intel(R) Xeon(R) Platinum 8358P CPU, 1T GB RAM, and a NVIDIA RTX 4090 24 GB GPU. In all scequences, the hyprparameters used in our experiments are listed in the supplementary materials.
@@ 4.1 Experiment Setup
Datasets. We conduct comparative experiments on widely used datasets, including 7 scenes from Mip-NeRF [1], 2 scenes from Tanks & Temples [27] and 2 scenes from Deep Blending [16]. All of these datasets do not provide precise geometric and semantic ground truth. To address this, we use the results obtained from the depth rendering pipeline in Normal-GS [61] as geometric priors for supervision. At the same time, we utilize open vocabulary queries to pre-process semantic labels with OpenSeed [75]. For accurate quantitative comparison experiments, we use the Replica [52] simulation dataset, which provides complete multimodal ground truth. We employ the pre-rendered Replica dataset from HiSplat [54], which includes 14 motion sequences across 7 scenes. The initial points in the Replica [52] dataset are obtained through mesh sampling, with a sampling interval of 5 in all sequences.
Baselines and Metrics. We compare our method with the existing SOTA NeRF-based dense framework Mip-NeRF [1], iNGP [44] and 3DGS framework [25], Scaffold-GS [40], GeoGaussian [33], Normal-GS [61], SemanticGaussian [15]. We evaluate the RGB rendering performance using PSNR, SSIM [60], and LPIPS [76]. The evaluation metrics for depth estimation we used are Abs.Rel. and RMSE. The evaluation metric we used for normal estimation is cosine similarity (CosSimi), for semantic segmentation, we employed mean Intersection over Union (mIoU).
@@ 4.2 Geometry Rendering Evaluation
We conduct a detailed quantitative comparison of depth and normal estimation against methods [25] [40] [22] [33] [61] on the Replica [52] dataset. The comparison results are shown in Fig. 3 and Tab. 1. Normal-GS [61] and our method optimize the surface normals, and both qualitatively and quantitatively outperform other 3DGS frameworks. Our approach also exhibits superior geometric accuracy in some extreme scenarios, particularly in texture-less areas, surpassing even Normal-GS [61]. More comparisons about geometry rendering can be found in supplementary materials.
@@
Figure 5: Quantitative comparison between our method and SGS-SLAM [32] in terms of RGB rendering and semantic segmentation clearly shows that our approach outperforms SGS-SLAM. The top-left corner of the image displays the PSNR value computed against the ground truth image.
Table 2: Quantitative geometric comparison between our method and other methods after sampling the reconstructed Gaussian model, compared to the initial points or mesh points. Best results are underlined.
@@
Method $|$ Metrics Radio Mean Std Hau-Dis Radio Mean Std Hau-Dis
Scaffold-GS [40] 0.0390 0.0299 0.0383 0.592 0.0476 0.0120 0.0083 0.141
Normal-GS [61] 0.0388 0.0303 0.0377 0.936 0.0477 0.0120 0.0082 0.131
Ours 0.0341 0.0272 0.0298 0.405 0.0339 0.0137 0.0081 0.078
We further compare the geometric impact after modifying our depth rendering method. Firstly, we complete the reconstruction process of a 3D Gaussian and then identify a detailed area in the reconstructed scene (as shown in  Fig. 4). It is evident that our method, compared to others, results in optimized Gaussian primitives that tend to be closer to the actual surface of the object. This is because our method integrates rotation and scale attribute into the optimization process, thereby achieving geometric consistency. Additionally, we perform point sampling [33] of the Gaussian primitives within specific small regions, followed by a Z-score statistical comparison with the initial ground truth and a Hausdorff distance comparison, as shown in Tab. 2. It is evident that our method achieves consistent geometric surface reconstruction. The calculation method using the Z-score statistic is explained in the supplementary materials.
Table 3: Quantitative comparison with the SGS-SLAM [32] method in terms of RGB rendering, semantic rendering and FPS. $T$ represents training time (minutes). Best results are underlined.
@@
Method $|$ Metrics $T$ PSNR FPS mIoU PSNR FPS mIoU PSNR FPS mIoU
SGS-SLAM [32] 212.6 36.28 96.4 0.976 39.14 72.2 0.978 36.94 112.3 0.944

### v2

@@ 1 Introduction
High-quality novel view synthesis (NVS) and fast multimodal 3D reconstruction remains a challenging problem in computer graphics and attracts widespread attention in fields such as autonomous driving [yan2024street, huang2025gaussianformer, huang2024gaussianformer, zheng2024gaussianad] and robotics [zhu20243d, lu2024manigaussian]. In recent years, neural radiance fields (NeRF) [verbin2022ref, barron2023zip, barron2022mip, Mildenhall20eccv_nerf] and 3D Gaussian Splatting (3DGS) [kerbl2024hierarchical, 3dgs] have made big breakthroughs in this field. NeRF uses coordinate-based networks to store 3D scene in multi-layer perceptrons (MLPs) via ray sampling, bringing in photo-realistic reconstruction. 3DGS, another approach, uses explicit 3D Gaussian primitives and modern GPU-accelerated splatting, and this design has been shown to enable high-quality real-time NVS by many 3DGS-based frameworks [scaffoldgs, lee2024compact, fan2024lightgaussian, niedermayr2024compressed, huang20242d, xie2024gs, ren2024octree, yan2024street, liu2024citygaussian, liu2024citygaussianv2]. However, the two types of existing methods still have many issues, which mainly focus on the following aspects. NeRF requires long training times, lacks easy scene editability, and its implicit 3D information storage method renders geometric reconstruction less intuitive [semanticnerf, Mildenhall20eccv_nerf]. Additionally, it performs poorly when applied to large-scale scenes [shi2025normal, blocknerf]. 3DGS and its related frameworks have more challenges: 1) Poor geometric structure and consistency: conventional image-centric 3DGS often ignores objects’ geometric structure and semantic info. It only uses Gaussian centers to represent depth, leaving out Gaussian rotation and scale attributes. This stops gradient propagation for geometric constraints (e.g. depth and normal), leading to inconsistent surface normals [scaffoldgs, 3dgs, li2024geogaussian, normalgs] and misalignment between Gaussians primitives and reconstructed surfaces. 2) Inefficient densification and pruning: some 3DGS methods create too many redundant Gaussians primitives to fit RGB images well [kerbl2024hierarchical, li2024geogaussian, 3dgs, yan2024street]. Though this improves metrics, it slows down rendering and uses more GPU memory, losing 3DGS’s real-time edge. 3) Poor unified multimodal framework: some 3DGS-related methods use separate frameworks for multimodal rendering [sgsslam, guo2024semantic], which causes repeated calculations and inconsistent 3D scene representations, failing to unify multimodal modeling. Existing studies have tried single or multiple scene modalities [scaffoldgs, normalgs, li2024geogaussian, deng2025omnimap], but there is still no unified framework that combines all these modalities while keeping storage efficient and rendering fast. This gap serves as the key motivation for our approach.
@@ 2 Related Works
Photo-Realistic Rendering and Densification. In the domain of photo-realistic reconstruction and rendering, NeRF and 3DGS represent implicit and explicit techniques, respectively. NeRF [Mildenhall20eccv_nerf, barron2022mip] and its variants pioneer the use of implicit neural representations for NVS, achieving high-fidelity rendering by optimizing a continuous volumetric scene network using multi-view images. However, this approach typically requires extremely long training times. Subsequent works incorporate geometric prior structures [yu2021plenoxels, barron2023zip, tang2024hisplat] or efficient encoding [muller2022instant] to accelerate the process. Despite these improvements, NeRF still faces inherent limitations, such as the difficulty in editing implicitly represented scenes. In contrast, 3DGS [3dgs, yang2024spectrally] employs explicit Gaussian primitives as scene representations. Through compression [scaffoldgs, lee2024compact, fan2024lightgaussian, niedermayr2024compressed, huang20242d] and efficient geometric structures [kerbl2024hierarchical, chen2024hac], several studies extend 3DGS to large-scale outdoor [xie2024gs, ren2024octree, yan2024street, li2025scenesplat] and urban [liu2024citygaussian, liu2024citygaussianv2] environments while maintaining high-quality reconstruction. But the aforementioned research primarily focuses on RGB-based rendering and seldom incorporates explicit geometric information of the environment. Moreover, impressive efforts have been made to enhance optimization and storage efficiency. SteepestGS [wang2025steepest] uncovers fundamental principles of density control. DashGaussian [chen2025dashgaussian] significantly accelerates 3DGS optimization. GeoTexDensifier [jiang2024geotexdensifier] utilizes geometric information for Gaussian densification, and Pixel-GS [zhang2024pixel] introduces pixel-level regularization for density control. These methods are regarded as post-processing techniques and lack integration within a unified framework.
@@
Geometry-Aware Rendering. Geometry-aware approaches enhance reconstruction accuracy by incorporating surfel [Huang20242DGS, normalgs, jiang2024gaussianshader, guedon2024sugar, dai2024high], normals [normalgs, shi2025normal], and relighting [gao2024relightable, wu2024deferredgs, ye20243d] capabilities, enabling more physically realistic renderings and interactions. Some studies [wang2021neus, oechsle2021unisurf, yariv2021volume, shi2025normal] attempt to achieve accurate geometry within the NeRF framework, and while significant progress has been made, these methods still suffer from inherent limitations that are difficult to overcome with implicit representations. 3DGS [3dgs, scaffoldgs] does not inherently excel in geometric reconstruction either, as its Gaussian primitives are independent of each other, making precise geometry recovery an open challenge. Several works unitize 3DGS or 2DGS to recover depth [3dgs, scaffoldgs, ververas2024sags, normalgs], surfel [Huang20242DGS, normalgs, jiang2024gaussianshader] and meshes [Huang20242DGS, guedon2024sugar, dai2024high], including normal estimation [normalgs] and relighting [gao2024relightable, wu2024deferredgs, ye20243d] techniques, which can effectively reconstruct smooth object surfaces. Nevertheless, these approaches remain largely limited to small-scale objects. When extended to large environments such as outdoor or urban scenes [loccoz20243dgrt, wu20253dgut], the required computational time remains substantial. Some methods attempt to incorporate SDF-based surfel constraints [yu2024gsdf, yariv2021volume, wang2023neus2, li2023neuralangelo, yariv2023bakedsdf], inverse rendering [shi2025gir, ye2024geosplating], or restrictions [li2024geogaussian, Keetha2023SplaTAMST] on Gaussian properties. However, these approaches lack rigorous analytical derivation for geometric modeling, employ relatively limited quantitative metrics, and demonstrate constrained generalizability across diverse scenarios. Consequently, the improvements achieve thus far remain marginal in terms of geometric fidelity.
@@
Semantic Rendering Frameworks. Several works attempt to integrate semantic information into both NeRF [semanticnerf, chou2024gsnerf] and 3DGS [guo2024semantic, qin2024langsplat], including techniques like Open-Vocabulary Segmentation [zhang2023simple, lai2025exploring, deng2025omnimap] and Language-Driven Semantic Segmentation [radford21a, liang2023open, luo2023segclip, li2022language], with promising results. Embedding semantic information facilitates downstream tasks such as scene understanding [jiang2025votesplat, zhou2024hugs], navigation [lei2025gaussnav, jin2024gs], and scene editing [ye2024gaussian, 3dgsdrag2025]. For instance, SGS-SLAM [sgsslam] directly uses semantic ground truth as supervision signal and rendering output in 3DGS, though it fails to formulate a unified 3DGS map representation. Feature-GS [zhou2024feature] extracts semantic features from the environment using 3DGS, and SEGS-SLAM [wen2025segsslam] enhances structural awareness with appearance embeddings. Nevertheless, such approaches exhibit limitations in reconstructing accurate environmental geometry and fail to achieve consistent reconstruction of scene geometry and semantics.
@@ 3.1 Unified Map Representation and Objectives
Objective. The objective of this framework is to render RGB color image, geometric depth, surface normals, semantic logits, and a gradient factor map (used in differentiable pruning) simultaneously within a unified rasterization pipeline. These rendered outputs are then compared against their corresponding 2D ground truth to compute respective loss, which guide the optimization of the 3D Gaussian scene through a differentiable backward process. To accelerate convergence and enhance computational efficiency, we derive the relevant analytical solutions and implement them using CUDA-accelerated gradient propagation. Finally, we can obtain a geometry consistent photo-realistic 3DGS map representation.
@@ 3.3 Differentiable Pruning
Pruning. After a fixed interval of iterations $K_{p}$ ($K_{p}=3000$ in our experiments), our framework directly eliminates Gaussians with abnormal gradient factor $\|k-\mathbf{1}\|_{1}>\mathcal{T}_{k}$ , effectively pruning the set and reducing the number of Gaussians to enhance overall efficiency. After pruning operation, the $k$ value of all Gaussians will be reset to 0.9 to maintain the sustainability of updates. Please refer to our code for details.
@@ 3.4 Multimodal Rendering Framework
Architecture Overview. Our system is built upon a modified Scaffold-GS [scaffoldgs] architecture. For each anchor point, we maintain a foundational 3D Gaussian representation (1). During every training iteration, multiple MLPs are used to predict attribute offsets for each anchor.
@@ 24   Save $T_{\text{final}}=T$;
Method $|$ Metrics Abs.Rel ↓ RMSE↓ CS ↑ Abs.Rel↓ RMSE↓ CS ↑ Abs.Rel↓ RMSE↓ CS ↑ Abs.Rel↓ RMSE ↓ CS↑
GeoGaussian [li2024geogaussian] 0.5828 0.8006 0.3046 2.8959 1.1092 0.1861 0.1336 0.3236 0.5700 0.1082 0.2592 0.5998
3DGS [3dgs] 1.2231 0.9183 0.2304 0.8711 0.6238 0.3201 0.2134 0.4092 0.4963 1.4422 1.0972 0.2905
Scaffold-GS [scaffoldgs] 2.2857 1.2102 0.2195 1.4609 0.9413 0.2854 4.0575 1.4786 0.1555 - - 0.1286
[…]
@@
Tile-based Forward Rasterization. We extend the original 3DGS’s tile-based rendering pipeline [3dgs]. During rasterization, for each pixel $p=(u,v)$, we calculate the direction $\mathbf{d}$ of the camera ray $r(t)$ and then compute the intersection $t_{1}/t_{2}$ with the Gaussian projected onto this pixel. Finally, we obtain the corresponding attributes using alpha blending rendering. The multimodal forward rasterization pipeline is outlined in Alg. 2. The corresponding backpropagation pipeline is described in the supplementary materials.
Loss Function. For the RGB image loss, we employ a combination of L1 loss $\mathcal{L}_{l1}$ and SSIM [wang2004image] loss $\mathcal{L}_{\text{ssim}}$. For the normal loss $\mathcal{L}_{\text{normal}}$, we use cosine similarity, which is calculated as
@@
Implementation Details. Our framework is implemented by CUDA using the LibTorch framework [libtorch], incorporating CUDA code for Gaussian Splatting and trained on a server with a 2.60 GHz Intel(R) Xeon(R) Platinum 8358P CPU, 1T GB RAM, and a NVIDIA RTX 4090 24 GB GPU. In all scequences, the hyprparameters used in our experiments are listed in the supplementary materials.
@@ 4.1 Experiment Setup
Datasets. We conduct comparative experiments on widely used datasets, including 7 scenes from Mip-NeRF [barron2022mip], 2 scenes from Tanks & Temples [knapitsch2017tanks] and 2 scenes from Deep Blending [hedman2018deep]. All of these datasets do not provide precise geometric and semantic ground truth. To address this, we use the results obtained from the depth rendering pipeline in Normal-GS [normalgs] as geometric priors for supervision. At the same time, we utilize open vocabulary queries to pre-process semantic labels with OpenSeed [zhang2023simple]. For accurate quantitative comparison experiments, we use the Replica [straub2019replica] simulation dataset, which provides complete multimodal ground truth. We employ the pre-rendered Replica dataset from HiSplat [tang2024hisplat], which includes 14 motion sequences across 7 scenes. The initial points in the Replica [straub2019replica] dataset are obtained through mesh sampling, with a sampling interval of 5 in all sequences.
Baselines and Metrics. We compare our method with the existing SOTA NeRF-based dense framework Mip-NeRF [barron2022mip], iNGP [muller2022instant] and 3DGS framework [3dgs], Scaffold-GS [scaffoldgs], GeoGaussian [li2024geogaussian], 2DGS [huang20242d], Normal-GS [normalgs], SemanticGaussian [guo2024semantic]. We evaluate the RGB rendering performance using PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable]. The evaluation metrics for depth estimation we used are Abs.Rel. and RMSE. The evaluation metric we used for normal estimation is cosine similarity (CosSimi), for semantic segmentation, we employed mean mIoU.
@@ 4.2 Geometry Rendering Evaluation
We conduct detailed qualitative and quantitative comparisons of depth and normal estimation against methods [3dgs] [scaffoldgs] [jiang2024gaussianshader] [li2024geogaussian] [normalgs] [huang20242d] on the Replica [straub2019replica] dataset. The comparison results are shown in Fig. 3 and Tab. 1. Normal-GS [normalgs] and our method optimize the surface normals, and both qualitatively and quantitatively outperform other 3DGS frameworks. Our approach also exhibits superior geometric accuracy in some extreme scenarios, particularly in texture-less areas, surpassing even Normal-GS [normalgs] and 2DGS [huang20242d]. More comparisons about geometry rendering can be found in supplementary materials.
@@
Figure 5: Quantitative comparison between our method and SGS-SLAM [sgsslam] in terms of RGB rendering and semantic segmentation clearly shows that our approach outperforms SGS-SLAM. The top-left corner of the image displays the PSNR value computed against the ground truth image.
Table 2: Quantitative geometric comparison between our method and other methods after sampling the reconstructed Gaussian model, compared to the initial points or mesh points.
@@
Method $|$ Metrics Radio↓ Mean↓ Std↓ Hau-Dis↓ Radio ↓ Mean↓ Std ↓ Hau-Dis ↓
[…]
2DGS [huang20242d] 0.0607 0.0984 0.1272 1.091 0.0628 0.0519 0.0244 0.462
[…]
We further compare the geometric impact after modifying our depth rendering method. Firstly, we complete the reconstruction process of a 3D Gaussian scene and then identify a detailed area in the reconstructed scene (as shown in  Fig. 4). It is evident that our method, compared to others, results in optimized Gaussian primitives that tend to be closer to the actual surface of the object. This is because our method integrates rotation and scale attribute into the optimization process, thereby achieving geometric consistency. Additionally, we perform point sampling [li2024geogaussian] of the Gaussian primitives within specific small regions, followed by a Z-score statistical comparison with the initial ground truth and a Hausdorff distance comparison, as shown in Tab. 2. It is evident that our method achieves consistent geometric surface reconstruction. The calculation method using the Z-score statistic is explained in the supplementary materials.
Table 3: Quantitative comparison with the SGS-SLAM [sgsslam] method in terms of RGB rendering, semantic rendering and FPS. $T$ represents training time (minutes). Best results are underlined.
@@
Method $|$ Metrics TT↓ PSNR ↑ FPS ↑ mIoU ↑ PSNR ↑ FPS ↑ mIoU↑ PSNR ↑ FPS ↑ mIoU ↑
SGS-SLAM [sgsslam] 212.6 36.28 96.4 0.976 39.14 72.2 0.978 36.94 112.3 0.944